A professional boxer hits his opponent with a 1160
Answer #1

F = 1160 N
t = 0.180 s

(a)
Impulse boxer imparts to his opponent = F * t
I = 1160 * 0.180
I = 208.8 kg m/s

(b)
F.t = m*v
208.8 = 112 * v
v = 1.86 m/s

(c)
Using Momentum conservation,
5 * Vrc = 112 * 1.86
Vrc = 41.6 m/s